BCAL Installation Schedule

From GlueXWiki
Jump to: navigation, search


  • BCAL schedule and activities for 2013
  • Reconstruction efforts


Attendees: Elton, David (JLab); George, Zisis, Shaun, Andrei (UofR)

Elton gave a detailed overview of the current and planned BCAL activities and timetables.

  1. 13 LGs have been glued so far. The newest glue batch had problems curing. A new batch arrived today and will be evaluated. Techs are trained for the gluing. Three persons are needed to do the job, including prep work. Shaun's trips to JLab alleviate the pressure on the techs, who can then turn to other tasks. The current gluing pace is 4 modules per week, so with 35 modules remaining we need 9 weeks more to finish. If all goes well, this should be completed by mid May.
  2. Mechanical process:
    1. Inspect LGs and glue them.
    2. Form the tedlar covers and tape them onto the LGs.
    3. Mount the LEDs. Inspect them visually (10 V, 100kHz rate).
    4. Move module to test station and use mini BCAL hardware. The final hardware is not quite ready. The alignment jig will arrive this week.
    5. Subassemblies (into the readout module). Copper cooling pipes are bent with a jig, fastened to the cooling plate with two brackets and the glued/soldered into place. The brackets stay on. Four thin termistor boards go on the cooling plate again in thermal contact.
    6. Other pres are needed but nobody has been assigned to them yet.
    7. Electrical guys will prepare boards.
    8. SiPMs sorted into bins and physically separated. 1000 units are still not at JLab, 360 that were returned to Japan from Chile were sent to JLab and left today for Chile. This delay may cause a problem in connection to the sorting and finding 40 units for each end.
  3. Magnet has delays. It is at 80 K, compressor fixed, one leak fixed, one small one still exists. Refrigerator is cooling down, no He yet. The plan is too cool it down, see if all ok and then warm it up to 80 K while the BCAL is being installed.
  4. BCAL modules can start being loaded into the cradle on the floor. This will relieve congestion on the Hall floor.
  5. Hardware will go on the ends of the modules before they go into the magnet's bore. If all goes well, in June we may be able to start with cosmics. There are questions on the schedule for the availability of the main electronics and racks. That schedule is being discussed this week. If not available, we will use the ESB setup in the hall as our DAQ.
  6. Regina crew should contribute productively in June and July.
  7. Contract end extension was discussed, specifically moving it from April 30 to May 31 or June 30. Elton will discuss it with management and procurement. Zisis will check with Danny for his availability.
  8. David was able to read Andrei's ROOT tree with the sampling fraction info. Andrei and Irina will work on neutron coefficients next (back burner) and then charged particles. For the latter, the entrance angle to the BCAL might be a better variable than the angle at the target.
  9. Collaboration will move to GEANT4 in April. This will not affect our BCAL reconstruction work, as it is mostly in mcsmear.
  10. We all agreed that reconstruction (for pi0's) is a priority. We will proceed by implementing a time-walk correction function with a single set of parameters for now, and then add parameters that depend on the energy deposition later (a la sampling fraction). We should plan code so that our temporary solution does not make it hard to implement the permanent one later on.
  11. David explained that we now have the three digitized quantities, energy and time from the FADC (good resolution on the time) and time from the TDC. Will and David will work on implementing these. Andrei will look at the code more globally, see if it is complete and investigate possible improvements such as energy deposition shapes as part of the algorithm, that might aid in photon-neutron separation.
  12. We agreed to meet every two weeks, on Thursdays at 11am JLab time with the first meeting on March 7th.
  13. Shaun will go to JLab March 12-29. Elton will cable up the second end of the mini BCAL setup trigger counter and Shaun will install a second trigger counter when he arrives.